# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ 5.2 Continuity
$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
Value $1\mp 2.155\;10^{-8}$ $1\pm 2.073\;10^{-8}$ $041\pm 6.065\;10^{-8} 1.1$$122\pm 1.328\;10^{-7}$
Uncertainty $0+2.127\;10^{-7}$ $0.201-1.358\;10^{-6}$ $0.407\pm 2.201\;10^{-7}$ $0.654\pm 2.784\;10^{-7}$
Table 1: The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$. The value and the uncertainty are expressed as the difference with the corresponding value and uncertainty of those of $(1\pm\delta x)^{n}$.
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$. Table 1 shows that the result of $(1\pm\delta x)^{n\pm d}$ where $0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former has an upper bound for $\delta x$, while the latter does not.
A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity. The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram. Figure 15 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
@@
In each case, the transition from $x=0$ to $x=0.2$ is continuous.
A statistical bounding range in variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative. An illustrative example is $\sin(x)$ shown in Figure 12.
A statistical bounding range in variance arithmetic cannot include any distributional zero because the result will diverge, such as at $x=0$ for both $(x\pm\delta x)^{c},c<1$ and $\log(x\pm\delta x)$.
@@ 6 Mathematical Library Functions
Figure 16: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $e^{x\pm\delta x}$, for different $x$ (x-axis), and different $\delta x$ (legend).
Figure 17: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\log(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
Figure 18: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $(1\pm\delta x)^{c}$, for different $c$ (x-axis), and different $\delta x$ (legend).
Figure 19: Error deviation for $(1\pm\delta x)^{c}$ as a function of $c$ and $\delta x$. The x-axis represents $c$ value between 2-2 and +3+3. The y-axis represents $\delta x$ value between $10^{-16}$ and $1$. The z-axis shows the corresponding error deviation.
Figure 20: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\sin(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
@@
Formula (2.20), (2.23), (2.26), and (2.29) are evaluated using the corresponding mathematical library functions exp, log, sin, and pow, respectively.
At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as:
  1. 1.
    Generate $10000$ samples from a Gaussian noise distribution, each with $\delta x$ as the distributional deviation, and construct $\tilde{x}$ by adding the sampled noise to $x$.
  2. 2.
    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between the outputs obtained using $\tilde{x}$ and $x$ as inputs.
  3. 3.
    Divide the value error by the result uncertainty to obtain the normalized error.
  4. 4.
    Calculate the standard deviation of the normalized errors to determine the error deviation.
6.1 Exponential
Figure 16 shows that the calculated uncertainties obtained using Formula (2.20) align closely with the measured value deviations for $e^{x+\delta x}$. Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
6.2 Logarithm
Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$. Figure 17 shows that the uncertainties calculated using Formula (2.23) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
6.3 Power
Figure 18 shows that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.29) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11. Figure 19 reveals that the error deviations of $\sin(x+\delta x)$ remain close to 11 for all input exponents $c$ and input uncertainties $\delta x>10^{15}$.
6.4 Sine
Figure 20 shows that the uncertainties calculated using Formula (2.26) correspond closely to the measured value deviations for $\sin(x+\delta x)$. It also reveals that $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$:
  • When $x=0$, $\sin(x)\simeq x$, so that $\delta^{2}\sin(x)\simeq(\delta x)^{2}$.
  • When $x=\pi/2$, $\sin(x)\simeq 1$, so that $\delta^{2}\sin(x)\simeq 0$.
Since $\sin(x)$ has distributional poles at $x=\pm\pi/2$, Figure 21 shows that the error deviation for $\sin(x+\delta x)$ equals $1$ except when $x=\pm\pi/2$ and $\delta x<10^{-8}$, matching the expected Delta distribution these poles. Elsewhere, the error deviations remain close to $1$.
6.5 Numerical Errors for Library Functions
The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$. Using either variance arithmetic or conventional floating-point library functions produces identical value errors. Figure 22 shows the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$. Figure 22 also shows that $e^{\log(x)}$ exhibits much smaller error than $\log(e^{x})$. For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise. The reason for the unexpectedly small value errors for $e^{\log(x)}-x$ is not yet clear.
The numerical error of the library function $x^{p}$ is computed as $(x^{p})^{1/p}-x$. Figure 23 shows that the normalized errors do not depend on either $x$ or $p$, resulting in an error deviation $0.548$.
The numerical errors of the library functions $\sin(x)$, $\cos(x)$, and $\tan(x)$ will be examined in greater detail in Section 9.
6.6 Summary
Formula (2.20), (2.23), (2.29), and (2.26) provide effective estimates of the corresponding library functions. When added noise exceeds $10^{-15}$ precision, ideal coverage is achieved except near a distributional pole where the error deviation approaches zero. In all other cases, proper coverage is attainable.
7 Matrix Calculations
7.1 Matrix Determinant
Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [14]. Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$14]. Formula (7.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [14]. The sub matrix $\mathbf{M_{i,j}}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (7.2) [14]. For discussion simplicity, sub determinant $|\mathbf{M}|_{i,j}$ in Formula (7.2) contains the permutation sign, which is different from the determinant of the sub matrix $|\mathbf{M}_{i,j}|$ [14] that treats the sub matrix $\mathbf{M_{i,j}}$ as an independent matrix. Formula (7.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [14].
$|\mathbf{M}|$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}x_{k,p_{k}};$ (7.1)
$|\mathbf{M}|_{i,j}$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{p_{i}=j}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\neq i}x_{k,p_{k}};$ (7.2)
$|\mathbf{M}|$ $=\sum_{j=1\dots n}|\mathbf{M}_{i,j}|x_{i,j}=\sum_{i=1\dots n}|\mathbf{M}|_{i,j}x_{i,j};$ (7.3)
Let $i_{1},i_{2},\dots>$} denote an ordered permutation of a subset from $1\dots n$, and $i_{1},i_{2},\dots]$} an unordered permutation [14]. Apply Formula (7.3) progressively to $M_{i,j}$, to expand Formula (7.2) as (7.4), and Formula (7.1) as (7.5). The (nm)(n-m)-by-(nm)(n-m) sub matrix in $|\mathbf{M}_{<i_{1}\dots i_{m}>_{n}[j_{1}\dots j_{m}]_{n}}|$ is obtained by deleting the rows in $\{i_{1}\dots i_{m}\}$ and the columns in $\{j_{1}\dots j_{m}\}$. This leads to Formula (7.6).
$|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{k\in\{i_{1}\dots i_{m}\}:p_{k}=j_{k}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\not\in\{i_{1}\dots i_{m}\}}x_{k,p_{k}};$ (7.4)
$|\mathbf{M}|$ $=\sum_{{[j_{1}\dots j_{m}]}_{n}}|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{k=1}^{m}x_{i_{k},j_{k}};$ (7.5)
$\left||\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\right|=\left||\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}|}\right|;$ (7.6)
[…]
$=\sum_{m=0\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$
$\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (7.8)
$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (7.9)
Formula (7.8) and (7.9) assume that the uncertainties of matrix elements are all independent of each other. This assumption maximized the result uncertainties. For discussion simplicity, other uncertainty assumptions are ignored in this paper.
7.2 Adjugate Matrix
Figure 24: Uncertainty means (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
The square matrix whose element is $a_{i,j}=(-1)^{i+j}|\mathbf{M}_{j,i}|$ is defined as the adjugate matrix [14] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$. Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [14]. Formula (7.10) and (7.11) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ [14].
$|\mathbf{M}|\;\mathbf{M}^{A}$ $=|\mathbf{M}^{A}|\;\mathbf{M};$ (7.11)
To test Formula (7.9):
  1. 1.
    A matrix $\mathbf{M}$ is constructed using random integers uniformly distributed in the range of $2^{8},+2^{8}]$, which has a distribution deviation of $2^{8}/\sqrt{3}$. The integer arithmetic ensures that $|\mathbf{M}|$, $\mathbf{M}^{A}$, and $|\mathbf{M}^{A}|$ are precise.
  2. 2.
[…]
The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test. As seen in Figure 24, the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$. Figure 25 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$. The existence of ideal coverage validates Formula (7.9).
Formula (7.10) defines the Forward Test. Figure 26 shows that the difference of the two sides of Formula (7.10) is precise zero, whether it is $\widetilde{\mathbf{M}}\times\widetilde{\mathbf{M}}^{A}-|\widetilde{\mathbf{M}}|\mathbf{I}$, or $\widetilde{\mathbf{M}}^{A}\times\widetilde{\mathbf{M}}-|\widetilde{\mathbf{M}}|\mathbf{I}$. The difference in Formula (7.10) is closer to Delta distribution for larger input precision, because without input noise, rounding errors dominate value errors. The validation of Formula (7.9) leads naturally to the validation of Formula (7.10).
Formula (7.11) defines the Roundtrip Test. Similarly, Figure 27 validates Formula (7.11). Because roundtrip test is no longer linear, error deviation in Figure 27 no longer increases linearly with input precision.
7.3 Floating Point Rounding Errors
Figure 28: Histograms of normalized errors of the adjugate matrix as a function of matrix size without input noise (legend).
Figure 29: Histograms of normalized errors of the adjugate matrix as a function of matrix size with $10^{-11}$ input noise (legend).
The significance of the conventional floating-point representation [8] has a $53$-bit resolution. As shown in Figure 28, the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 28. The rounding error is also the reason why only at matrix size 8 and input noise precision $10^{-10}$, error deviation is no longer $1$ in Figure 25.
With $10^{-11}$ noises added to the input, Figure 29, the distribution becomes Gaussian with a hint of Delta distribution. Such Delta-like distribution persists until the input noise precision reaches $10^{-10}$, which is also the transition of the two trends in Figure 24. Figure 29 shows the distribution when the input noise precision is $10^{-11}$, where the distinction due to matrix size vanishes.
7.4 First Order Approximation
[…]
$|\widetilde{\mathbf{M}}|\simeq\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}(x_{k,p_{k}}+\tilde{x}_{k,p_{k}});\;\;\;\Rightarrow\;\;\;\delta^{2}|\mathbf{M}|\simeq\sum_{i}^{n}\sum_{j}^{n}M_{i,j}(\delta x_{i,j})^{2};$ (7.12)
Figure 30 contains the result of applying Formula (7.12). It is very similar to Figure 25, validating Formula (7.12).
7.5 Matrix Inversion
$\mathbf{M}^{-1}\equiv$ $\;\mathbf{M}^{A}/|\mathbf{M};$ (7.13)
$\mathbf{M}^{-1}\times\mathbf{M}=\mathbf{M}\times\mathbf{M}^{-1}=\mathbf{I};$ (7.14)
$(\mathbf{M}^{-1})^{-1}=\mathbf{M};$ (7.15)
An inverse matrix is defined by Formula (7.13), which satisfies Formula (7.14) and (7.15) [14]. However, this definition is seldom used conventionally to calculate inverse matrix due to large uncertainty response ratio which appears as small input uncertainties causing large output uncertainties [12][14]. Traditionally, matrix condition number [14] is a proxy for uncertainty response ratio of matrix inversion. In Formula (7.13), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$. Figure 31 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices. As a reference, Figure 31 presents the Hilbert matrix [14] for each matrix size, and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number. Thus, determinant precision can replace condition number to estimate uncertainty response ratio of matrix inversion.
$\mathbf{M}=$ $\;\left(\begin{matrix}w,x\\ y,z\end{matrix}\right);\;\;\;\mathbf{M}^{-1}=\frac{\left(\begin{matrix}z,-x\\ -y,w\end{matrix}\right)}{wz-xy};$ (7.16)
$\delta^{2}\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}z^{4},x^{2}z^{2}\\ y^{2}z^{2},x^{2}y^{2}\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}y^{2}z^{2},w^{2}z^{2}\\ y^{4},w^{2}y^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{4}}+$ (7.17)
$\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}2z^{3},-2xz^{2}\\ -2yz^{2},2xyz\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}2y^{2}z,-2wyz\\ -2y^{3},2wy^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{3}}+$ (7.18)
$\frac{\left(\begin{matrix}2x^{2}z,-2x^{3}\\ -2wxz,2wx^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}2wxy,-2w^{2}x\\ -2w^{2}y,2w^{3}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{3}};$
Variance arithmetic is path-independent, so it computes inverse matrix directly from the definition of Formula (7.13). For example, for the inverse matrix $\mathbf{M}^{-1}$ of size $2$ in Formula (7.16), Formula (7.17) and (7.18) present first-order approximations for variance $\delta^{2}\mathbf{M}^{-1}$, and bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, respectively. The resulting uncertainties of determinant $\delta^{2}|\mathbf{M}|$ and inversion $\delta^{2}\mathbf{M}^{-1}$, as well as the resulting bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, are not linear to each input uncertainty $\delta x_{i,j}$. Inverse variance $\delta^{2}\mathbf{M}^{-1}$ contains sum of all input variance $(\delta x_{i,j})^{2}$ and their higher-order permutation products, such that the uncertainty response ratio for matrix inversion should roughly equal the matrix size. Thus, an uncertainty response ratio for matrix inversion is inherently large, independent of computational path. A seemly small uncertainty response ratio using Gaussian elimination [12][14] is probably a path-dependent artifact; while the conventionally “bad” method of applying Formula (7.13) for matrix inversion [12][14] is actually the correct one, except missing resulting uncertainty when using floating-point arithmetic. All conventional path-dependent results are questionable. Only statistical Taylor expansion presents a complete picture of resulting uncertainty for an analytic expression.
It is doubtful if Formula (7.15) still holds for uncertainty in statistical Taylor expansion, because it seems that uncertainty response ratio can only increase in matrix inversion according to Formula (7.17). On the other hand, because the bias in Formula (7.18) can be either positive or negative, it is expected that Formula (7.15) still holds for value.

### v17

@@ 5.2 Continuity
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number $n$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
5.3 Distributional Hole
@@
Unlike a distributional zero, a statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$. The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram. Figure 13 illustrates the error distributions of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
@@
In each case, the transition from $x=0$ to $x=0.2$ is continuous for the resulting mean and deviation.
6 Matrix Calculations
$\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (6.1)
$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (6.2)
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Formula (6.1) and (6.2) gives the statistical Taylor expansion of the determinant of a square matrix $\mathbf{M}$ of size $n$, when the uncertainties of matrix elements are all independent of each other. In Formula (6.2), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices.
The square matrix whose element is $(-1)^{i+j}|m_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$, in which $m_{j,i}$ is the element of $\mathbf{M}$ at row index $j$ and column index $i$. Each $m_{j,i}$ is an integer between $2^{8},+2^{8}]$ so that $\mathbf{M}^{A}$ can be calculated precisely using integer arithmetic. By adding Gaussian noise of deviation $\delta x$ to each $m_{j,i}$, $\widetilde{\mathbf{M}}$ is created. The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, while the result deviation is calculated by Formula (6.2), so that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test. Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$. This abnormality is caused by floating-point rounding errors only at this point, when the element values exceed the significand range of 64-bit floating-point representation, to result in a error deviation of $3.2$. Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.
$\mathbf{M}^{-1}$ $\equiv\;\mathbf{M}^{A}/|\mathbf{M}|;$ (6.4)
$\delta^{2}\left(\begin{matrix}w,x\\ y,z\end{matrix}\right)^{-1}\simeq$ $\;\frac{\left(\begin{matrix}z^{4},x^{2}z^{2}\\ y^{2}z^{2},x^{2}y^{2}\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}y^{2}z^{2},w^{2}z^{2}\\ y^{4},w^{2}y^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{4}}+$ (6.5)
Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [20]. Formula (6.3) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.4) [20]. $\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value so that it is a value test. Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the value should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate. The matrix size is the specific dimension in Figure 15.
Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.4), its result variance is very complicated using Formula (2.3). For example, Formula (6.5) shows the simplest case for Formula (6.4): the first-order approximation of a 2x2 matrix. Contrary to traditional approach, statistical Taylor expansion uses Formula (6.4) for matrix inversion instead of Gaussian elimination [5], because logically, the result should be symmetric for all matrix elements, as what Formula (6.5) has demonstrated.
In Formula (6.4), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$. Figure 16 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices. As a reference, Figure 16 presents the Hilbert matrix [20] for each matrix size and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number. Thus, determinant precision can replace matrix condition number.
@@ 7 Mathematical Library Functions
Basic Function $e^{x\pm\delta x}$ $\log(x\pm\delta x)$ $(1\pm\delta x)^{c}$ $\sin(x\pm\delta x)$
Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.978\pm 0.150$
Table 1: The result error deviation of selected basic functions using variance arithmetic when $\delta x>10^{-15}$. The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverge regions are excluded. The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
@@
Table 1 shows that Formula (2.15), (2.17), (2.20), and (2.22) provide almost perfect deviation for the result values which are calculated using corresponding math library functions by sampling from $x\pm\delta x$ in which $\delta x$ is Gaussian input noise deviation. Table 1 shows the results from Python, which differs slightly from those from either C++ or Java.
The result error deviation is similar to Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$. The coverage is only proper when $\delta x<10^{-15}$.
Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$. Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation. As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$. The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ will be examined in greater detail in Section 10.
To test $f^{-1}(f(x))-x=0$ when $\delta x=0$ for the library functions:
  • Figure 18 shows that the value errors in $e^{\log(x)}-x$ are much less than those in $\log(e^{x})-x$. For $\log(e^{x})-x$, the error deviation is $0.41$ when $|x|\leq 1$, or 0 otherwise.
  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, independent on neither $x$ nor $p$.
The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has almost no value error, are not clear.